A source of light is located at corner inside a 3m deep pool. If you were standing 4m away on the side of the pool.
(1) will the light be totally reflected?
(2) if you were standing directly above the source of light ,what will be the apparent depth of the light?
1
Expert's answer
2020-02-19T13:24:40-0500

(1)


From the above figure "sin \theta=\dfrac{4}{\sqrt{4^2+3^2}}=\dfrac{4}{5}"


The angle of incidence here is "sin^{-1}4\/5 = 53^0"

For the water-air boundary, the critical angle is 48.6-degrees

So when angle of incidence greater than critical angle then total internal reflection takes place

here total internal reflection will take place


(2). "Apparent\ depth = \dfrac{real \ depth} {μ} =\dfrac 3{(4/3)} = \dfrac 94"